# Multi-Agent Reinforcement Learning from Human Feedback: Data Coverage and Algorithmic Techniques

### v2

@@ 6.1 Environments
Our experiments involved 3 Multi-Agent Particle Environments (MPE) implemented with JaxMARL codebase (Rutherford et al., 2023). Specifically, we utilized the Spread, Tag, and Reference scenarios, which covers three different levels of communication requirements. Spread-v3 contains a group of agents and target landmarks, where the objective is to cover as many landmarks as possible while avoiding collisions. Tag-v3 contains two opposing groups, where quicker "preys" need to escape from "predators". To ensure a fair comparison of different predator cooperation policies, we fixed a pretrained prey agent. Reference-v3 involves two agents and potential landmarks, where the agents need to find each one’s target landmark to receive a high reward. The target landmark of each agent is only known by the other agent at first, necessitating communication before they can move toward the correct landmark. A more detailed description of the tasks and their associated challenges is provided in Appendix B.2.
@@ 6.2 The Importance of Dataset Diversity
  • Expert policy: $n$ expert agents. Trained with VDN till convergence.
  • Rookie policy: $n$ rookie agents. Trained with VDN with early stop.
@@
The experimental results are presented in Table 2. We can clearly observe that trivial comparisons (Mix-Expert vs. Pure-Expert) play a crucial role in learning an appropriate reward. A dataset containing only expert policies proves hard for the reward model to learn from the very beginning. In more challenging tasks, as reflected by higher MSE, the importance of unilateral coverage and diversity becomes more pronounced. Suboptimal cooperative agent actions assist the reward model in identifying the Nash equilibrium, thus highlighting key reward signals more effectively, as suggested by Theorem 2. Additionally, we found that in a diversified dataset, the variance across all training runs was lower compared to other datasets.
Beyond test returns, we also visualized the reward predictions and utilized the MSE between standardized reward predictions and ground truth as a measure of prediction quality. Although this MSE is not necessarily directly correlated with final performance—since accurate predictions do not need to match the shape of the ground truth—it can serve as a useful empirical metric.
@@
Table 1: Final datasets mixed with various ratios. The overall dataset size is kept to 38400 trajectories. More details list in Appendix B.1.
Spread-v3 Tag-v3 Reference-v3 Spread-v3 Tag-v3 Reference-v3
Return Return Return MSE MSE MSE
Diversified -21.63 $\pm$ 0.51 28.94 $\pm$ 1.04 -19.59 $\pm$ 0.55 0.51 1.04 0.55
Mix-Unilateral -20.98 $\pm$ 0.56 32.04 $\pm$ 1.12 -20.60 $\pm$ 1.15 0.56 1.12 1.15
Mix-Expert -21.11 $\pm$ 1.16 33.01 $\pm$ 0.70 -20.38 $\pm$ 1.06 1.16 0.70 1.06
Pure-Expert -21.45 $\pm$ 1.01 40.14 $\pm$ 0.75 -28.19 $\pm$ 0.92 1.01 0.75 0.92
Table 2: Test returns and the mean squared error (MSE) between the standardized predicted rewards and the standardized ground truth rewards. Detailed descriptions of the datasets are provided in Section 6.2. In environments with smoother reward distributions, such as Spread-v3 and Reference-v3, dataset diversity primarily reduces variances. However, in more challenging environments, such as Tag-v3, dataset diversity plays a substantially more significant role.
6.3 Experiments for Reward Regularization
In Figure 2, we examined the effectiveness of our proposed reward regularization technique. Figure 2(a) demonstrates that without regularization, the learned rewards tend to be sparse and spiky compared to the ground truth rewards. The predicted rewards in (b1) and (b3) were learned using our reward regularization, while those in (b2) were learned without it ($\alpha=0$). Our results indicate that the reward regularization technique produces much smoother reward functions. We also quantitatively compare learned rewards with and without using reward regularization in finding the policy (cf. the first and the last column of Table 3).
We also observe that the rewards often exhibit temporal continuity, which can create greater discrepancies with the sparse, pulse-like ground truth. Notably, we found that adding stronger regularization does not necessarily lead to underfitting of the reward model; in some cases, it even helps the model converge to a lower training loss. Detailed parameters and experimental results are provided in the appendix (cf. Table 7). We attribute this to the role of regularization in preventing the model from overly relying on shortcuts.
@@ 6.3 Experiments for Reward Regularization
Table 3: Comparison of test return per agent with different parameters. Standard pipeline take KL reward coefficient $\beta=1$ for Spread-v3, Reference-v3 and $\beta=10$ for Tag-v3, and the MSE reward regularization coefficient $\alpha$ is set to the optimal value for fixed $\beta$. All the agents are trained on Diversified Dataset across 5 random seeds. This experiment shows the importance using reward regularization and reference policy.
6.4 Experiments for Imitation Learning
Intuitively, the choice of the KL coefficient $\beta$ depends on the quality of the reward model $\phi$ and the dataset. As shown in Table 3, simpler environments (e.g., Spread and Reference) require less KL restriction, while harder environments (e.g., Tag) need stronger restriction to avoid suffering from the extrapolation problem. A large $\beta$ can amplify the approximation error of imitation learning, potentially overshadowing the reward signal, whereas a small $\beta$ may fail to introduce enough pessimism to effectively mitigate extrapolation errors. In environments with relatively simple reward structures, such as Spread-v3 and Reference-v3, the standard choice of $\beta=1$ is near optimal. However, in more complex scenarios where the reward is harder to derive, increasing $\beta$ to 3 yields better performance.
@@ 7 Discussion
In this paper, we proposed dedicated algorithmic techniques for offline MARLHF and provided theoretical justification for the unilateral dataset coverage condition. We believe our work is a significant step towards systematically studying MARLHF and offers a foundational framework for future research in this area. The flexibility of our framework allows for application across a wide range of general games, and our empirical results validate the effectiveness of our proposed methods in various scenarios.
Looking ahead, there is significant potential to extend this work to more complex, real-world scenarios, particularly by integrating Large Language Models (LLMs) into multi-agent systems. Future research will focus on fine-tuning and aligning LLMs within MARLHF, addressing challenges such as increased complexity and the design of effective reward structures.
@@ A.2 Unilateral Policy Coverage
We use MLE for reward estimation for each player. For simplicity, we omit the subscript $i$ for player $i$. Note that the reward function can be expressed as $r_{\theta}(\tau)=\sum_{h=1}^{H}r_{h}(s_{h},\mathbf{a}_{h})=\left\langle\psi(\tau),\theta\right\rangle$, where $\theta=[\theta_{1},\theta_{2},\cdots,\theta_{H}]$. At each step $h$, we minimize the NLL loss:
@@
By Lemma 5, the confidence region is
@@
We define the optimistic reward and the pessimistic reward as
@@
By Lemma 5, with probability at least $1-\delta$, we have $\theta\in\Theta$. Thus we have
@@
On the other hand, we have
@@
Similarly, we have
@@
With probability at least $1-\delta$, we have
$\underline{V}^{\pi}_{h,i}(s,\mathbf{a})\leq V^{\pi}_{h,i}(s,\mathbf{a}),\overline{V}^{\dagger,\pi_{-i}}_{h}(s,\mathbf{a})\geq V^{\dagger,\pi_{-i}}_{h}(s,\mathbf{a}),\forall h\in[H],i\in[m],s\in\mathcal{S},\mathbf{a}\in\mathcal{A}.$
@@
We prove the arguments by mathematical induction. The arguments hold for step $H+1$ as all quantities are 0. Suppose step $h+1$ holds and we consider step $h$. For the first argument, we have
@@
For the second argument, we have
@@
By Lemma 6, we have
@@
Similarly, we have
@@
The proof is similar to the proof for Lemma 21 in [Cui and Du, 2022b]. ∎
@@ B.1 Implementation Details
The models used in our experiments are designed to effectively handle the complexities of multi-agent environments. Our reward model $r_{\phi}$ is a fully connected neural network, featuring action and observation embedding layers followed by hidden layers. The MADPO agent network uses RNN to output its Q-values, enabling the agent to make informed decisions based on its observations and learned policies. Table 5 lists the main hyperparameters used in our experiments, while other details can be checked in our codebase: https://github.com/NataliaZhang/marlhf.
@@
MSE Loss Coefficient 1
Reference Log Coefficient 1
@@
Number of Environment Timesteps 130
Number of Environments 128
Reward Model Type MLP (Spread, Reference), RNN (Tag)
@@
MLP Layer Dimension 128 (Spread), 64 (Reference)
@@
VDN Learning Rate 1e-3
VDN Epochs 1e4 (Spread, Reference), 1e6 (Tag)
VDN Batch Size 64
@@
As mentioned in Section 6.2, we collected trajectories in different environments using various policies to ensure a diverse dataset. Our dataset size contains 38400 trajectories in each environment. The number of trajectory pairs is chosen as 10 times the number of trajectories. Preference tags were then generated for these trajectory pairs in the mixed datasets. To adjust the randomness of the preferences, a steepness parameter was introduced as a scalar of the standardized reward. This configuration ensures a comprehensive dataset that can effectively support the evaluation of our methods.
@@ B.2 Tasks Descriptions
    • Challenge: The adversaries only gets reward at the timestep of catching a good agent, so recovering the reward distribution across time becomes a challenging work. Note that the original environment is a 1v3 adversary game, and we convert it into a 3-agent cooperative game for better evaluation by fixing the good agent with a MAPPO pretrained policy.

### v3

@@ 6.1 Environments
Our experiments involved three Multi-Agent Particle Environments (MPE), including Spread-v3, Tag-v3 and Reference-v3, and Overcooked environment implemented with JaxMARL codebase (Rutherford et al., 2023). Spread-v3 contains a group of agents and target landmarks, where the objective is to cover as many landmarks as possible while avoiding collisions. Tag-v3 contains two opposing groups, where quicker "preys" need to escape from "predators". To ensure a fair comparison of different predator cooperation policies, we fixed a pretrained prey agent. Reference-v3 involves two agents and three potential landmarks, where the agents need to find each one’s target landmark to receive a high reward. The target landmark of each agent is only known by the other agent at first. Overcooked involves two agents moving and operating objects in a gridworld. A more detailed description of the tasks and their associated challenges is provided in Appendix B.2.
@@ 6.2 The Importance of Dataset Diversity
  • Expert policy: $n$ expert agents. Trained with online RL algorithms till convergence.
  • Rookie policy: $n$ rookie agents. Trained with online RL algorithms with early stop.
@@
Due to the inherent limitations of offline reinforcement learning (RL) in action selection dictated by the dataset, the effectiveness of learning is often strongly correlated with dataset quality, i.e. the level of expertise demonstrated in the dataset. However, the results in preference-based MARL experiments partially diverge from this conventional conclusion. While the quality of the dataset remains critical, experiments on Reference-v3 and Overcooked (Table 2) indicate that diversity and unilateral data can significantly enhance the performance of the reward model, thereby facilitating learning.
The main experimental results are presented in Table 2 and Table 3.Among all the experiments, apart from the experiments on Tag-v3, where the high operational precision requirements make data quality more critical than diversity, the other three environments validate our conclusions across all algorithms.
@@
Table 1: Final datasets mixed with various ratios. The overall dataset size is kept to 38400 trajectories for MPE, and 960 trajectories for Overcooked. (cf. Appendix B.1)
Algorithm Dataset Spread-v3 Tag-v3 Reference-v3 Overcooked
VDN with Diversified -21.16 $\pm$ 0.54 29.28 $\pm$ 1.08 -18.89 $\pm$ 0.60 238.89 $\pm$ 3.50
Pessimism Penalty Mix-Unilateral -21.03 $\pm$ 0.44 36.65 $\pm$ 0.70 -18.80 $\pm$ 0.63 221.80 $\pm$ 26.66
Mix-Expert -20.98 $\pm$ 0.54 35.96 $\pm$ 0.86 -18.80 $\pm$ 0.44 35.26 $\pm$ 55.19
Pure-Expert -21.01 $\pm$ 0.57 39.55 $\pm$ 0.77 -28.97 $\pm$ 2.89 3.36 $\pm$ 7.19
Table 2: In the simplest environment, Spread-v3, different dataset gives similar performance. In Tag-v3 environment, where precise actions are required, the quality of the dataset (proportion of expert demonstration) is more important than diversity. In contrast, in Overcooked environment, which focuses on strategy learning and demands less on precision, dataset diversity contributes to improved stability, with Unilateral playing a particularly critical role. In the Reference-v3 environment, which balances the need for precision and strategic, the importance of both factors is more balanced, but non-expert data is still necessary.
@@
Algorithm Dataset Spread-v3 Reference-v3 Overcooked
MAIQL Diversified -25.33 $\pm$ 1.40 -22.15 $\pm$ 0.55 16.59 $\pm$ 11.22
Mix-Unilateral -23.25 $\pm$ 1.06 -23.22 $\pm$ 1.37 0.00 $\pm$ 0.00
Mix-Expert -23.26 $\pm$ 0.90 -24.21 $\pm$ 1.60 0.00 $\pm$ 0.00
Pure-Expert -26.01 $\pm$ 1.53 -29.47 $\pm$ 1.65 0.00 $\pm$ 0.00
MABCQ Diversified -20.02 $\pm$ 0.64 -17.64 $\pm$ 0.43 239.34 $\pm$ 1.67
Mix-Unilateral -19.47 $\pm$ 0.33 -17.64 $\pm$ 1.11 215.01 $\pm$ 65.43
Mix-Expert -19.42 $\pm$ 0.17 -17.88 $\pm$ 0.78 50.32 $\pm$ 82.82
Pure-Expert -20.56 $\pm$ 0.38 -25.90 $\pm$ 1.11 1.14 $\pm$ 3.46
Table 3: Test returns of MAIQL and MABCQ. In the experimental results, we can observe a clear preference toward more diversified datasets. Compared to our method and BCQ, which directly calculate $\max_{a}Q$ for Bellman updates, IQL employs expectile regression to estimate it. So MAIQL demands higher accuracy of the reward model. Consequently, the performance improvements brought by dataset diversity are also more pronounced in MAIQL experiments.
@@
Overcooked 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 149.53 $\pm$ 86.74 238.89 $\pm$ 3.50 240 $\pm$ 0.00 240 $\pm$ 0.00
Table 4: Comparison of test return with different hyperparameters. Standard pipeline take pessimism coefficient $\beta=1$ for Spread-v3, Reference-v3 and $\beta=10$ for Tag-v3, Overcooked, and the MSE reward regularization coefficient $\alpha$ is set to the optimal value for fixed $\beta$. All the agents are trained on Diversified Dataset across 10 random seeds. Results show that larger $\beta$ always gives better performance and a proper positive $\alpha$ can improve performance.
6.3 Other Ablation Studies
Reward regularization
In Figure 2, we examined the effectiveness of our proposed reward regularization technique. Figure 2(a) demonstrates that without regularization, the learned rewards tend to be sparse and spiky compared to the ground truth rewards. We also observe that the rewards often exhibit temporal continuity, which can create greater discrepancies with the sparse, pulse-like ground truth. Notably, we found that adding stronger regularization does not necessarily lead to underfitting of the reward model; in some cases, it even helps the model converge to a lower training loss. Detailed parameters and experimental results are provided in the appendix (cf. Table 8). We attribute this to the role of regularization in preventing the model from overly relying on shortcuts.
Pessimism coefficient
Due to the clipping in Equation 3, excessively large $\beta$ values will not dominate the entire reward function. As a result, larger $\beta$ values almost never degrade the agent’s performance in our experiments (Table 4). This allows us to increase $\beta$ with relative confidence. Therefore, we generally recommend setting $\beta$ to a value between 10 and 100 for optimal performances.
Scalability
We also tested the scalability on Spread-v3. While our current approach manages the scaling of agents without introducing new problems, it does not specifically address the inherent issues of instability and complexity that are well-documented in traditional MARL (cf. Appendix B.3).
@@ 7 Discussion
In this paper, we proposed dedicated algorithmic techniques for offline PbMARL and provided theoretical justification for the unilateral dataset coverage condition. We believe our work is a significant step towards systematically studying PbMARL and offers a foundational framework for future research in this area. The flexibility of our framework allows for application across a wide range of general games, and our empirical results validate the effectiveness of our proposed methods in various scenarios.
Looking ahead, there is significant potential to extend this work to more complex, real-world scenarios, particularly by integrating Large Language Models (LLMs) into multi-agent systems. Future research will focus on fine-tuning and aligning LLMs within PbMARL, addressing challenges such as increased complexity and the design of effective reward structures.
@@ A.2 Unilateral Policy Coverage
In our framework, we utilize Maximum Likelihood Estimation (MLE) to estimate the reward function for each player. For simplicity, we omit the subscript $i$ for player $i$. Note that the reward function can be expressed as $r_{\theta}(\tau)=\sum_{h=1}^{H}r_{h}(s_{h},\mathbf{a}_{h})=\left\langle\psi(\tau),\theta\right\rangle$, where $\theta=[\theta_{1},\theta_{2},\cdots,\theta_{H}]$ represents the parameters we aim to optimize. At each step $h$, we minimize the NLL loss:
@@
This optimization problem helps in learning a reward function that aligns well with the observed data.
By Lemma 5, we establish a confidence region for the estimated parameters $\theta$:
@@
We define the optimistic reward and the pessimistic reward as follows:
@@
This lemma establishes bounds on the reward function $r_{h,i}(s,\mathbf{a})$. Specifically, the optimistic reward and pessimistic reward are under constrains at a high probability, up to a margin of error determined by $C_{r}$ and the norms of the feature representation $\overline{\psi}_{h}(s,\mathbf{a})$.
@@
We begin by referencing Lemma 5, which establishes that with probability at least $1-\delta$, the estimated parameters $\theta$ reside within the confidence region $\Theta$. This allows us to assert the following relationships:
@@
Next, we quantify the deviation between the optimistic reward and the true reward:
@@
A similar argument can be applied to show that the pessimistic reward is also bounded:
@@
With probability at least $1-\delta$, the following bounds hold for all time steps $h\in[H]$, for each agent $i\in[m]$, for all states $s\in\mathcal{S}$, and actions $a\in\mathcal{A}$:
$\underline{V}^{\pi}_{h,i}(s,a)$ $\leq V^{\pi}_{h,i}(s,a),$
$\overline{V}^{\dagger,\pi_{-i}}_{h}(s,a)$ $\geq V^{\dagger,\pi_{-i}}_{h}(s,a).$
@@
We prove the statements by mathematical induction. The base case for step $H+1$ holds trivially, as all quantities are zero at this final step. Suppose step $h+1$ holds and we consider step $h$. We will show they also hold for step $h$. For the first argument of pessimistic value, we have
@@
For the second argument of optimistic value, we have
@@
These complete the induction step and shows the lemma holds for all steps. ∎
@@
This lemma establishes bounds on the difference between the expected value of the policy and its lower or upper estimates. Using properties of the Bellman operator, we can show that the difference between $V_{1,i}^{\pi}(s_{1})$ and $\underline{V}_{1,i}^{\pi}(s_{1})$ is controlled by the expected norms of the feature representations $\psi$ and $\overline{\psi}$, scaled by constants $C_{\mathcal{P}}$ and $C_{r}$.
@@
We prove each bound separately, utilizing the results from Lemma 6.
For the first bound of pessimistic value estimation, we analyze the expected action-value $Q^{\pi}_{1,i}(s_{1},\mathbf{a})$ and its lower estimate, applying the Bellman operator iteratively. We have
@@
Similarly, for optimistic value estimation, we have
@@
[…]
$\leq\sum_{i\in[m]}\left[\overline{V}_{1,i}^{\dagger,\pi^{*}_{-i}}(s_{1})-\underline{V}_{1,i}^{\pi}(s)\right].$
Utilizing Lemma 2, it is straightforward to derive this lemma. The proof is similar to the proof for Lemma 21 in [Cui and Du, 2022b]. ∎
@@ Appendix B Experiment Details
Algorithm 4 Pipeline of Preference-Based Multi-agent Reinforcement Learning
1:Input: Dataset $\mathcal{D}=\{\tau_{i},\tau_{i}^{\prime},\mathbf{y}_{i}\}_{i=1}^{N}$.
2:Train a agent-wise reward model $r_{\phi}$;
3:Train a imitation model $\pi_{b}$;
4:Apply MARL algorithm to learn $\pi_{\mathbf{w}}$ with distribution-based pessimism from $\pi_{b}$;
5:return $\pi_{\mathbf{w}}$.
@@ B.1 Implementation Details
Pipeline
Our pipeline for Preference-Based Multi-Agent Reinforcement Learning (PbMARL), detailed in Algorithm 4, outlines the key steps for training agents using preference datasets. The process begins with training an agent-specific reward model $r_{\phi}$, followed by learning a imitation model $\pi_{b}$. The final policy $\pi_{\mathbf{w}}$ is then optimized using a MARL algorithm with distribution-based pessimism derived from $\pi_{b}$.
@@
The models used in our experiments are designed to effectively handle the complexities of multi-agent environments. Our reward model $r_{\phi}$ is a fully connected neural network, featuring action and observation embedding layers followed by hidden layers. The MADPO agent network uses RNN to output its Q-values, enabling the agent to make informed decisions based on its observations and learned policies. MABCQ and MAIQL, as tested in Section  6, are modified versions of BCQ Fujimoto et al. [2019] and IQL Kostrikov et al. [2021] tailored for MARL. Similar to VDN, the Q-functions in MABCQ and MAIQL are designed to represent the cumulative rewards of all agents collectively. This adaptation ensures compatibility with the multi-agent setting while preserving the theoretical and practical foundations of the original algorithms. In MAIQL, the coefficient for expectile regression, $\tau$, is consistently set to 0.95 across all experiments. For MABCQ, random noise addition is omitted due to the discrete action spaces in the tested environments, and the VAE is replaced with a policy generator trained via imitation learning. Table 6 lists the main hyperparameters used in our experiments, while other details can be checked in our codebase: https://github.com/NataliaZhang/marlhf.
@@
MSE Loss Coefficient $\alpha$ 1 (MPE), 1e-3 (Overcooked)
Pessimism Coefficient $\beta$ 1 (Spread, Reference), 10 (Tag, Overcooked)
@@
Episode length 26 (MPE), 400 (Overcooked)
Reward Model Type MLP (Spread, Reference, Overcooked), RNN (Tag)
@@
MLP Layer Dimension 64
@@
Policy Model Learning Rate 1e-3
Policy Model Epochs 1e4 (Spread, Reference), 1e6 (Tag), 1e5 (Overcooked)
Policy Model Batch Size 128
MAIQL Expectile Regression Coefficient 0.95
Optimizer Adam
@@
As mentioned in Section 6.2, we collected trajectories in different environments using various policies to ensure a diverse dataset. Each dataset contains 38400 trajectories in each MPE environment and 960 trajectories in Overcooked. The number of trajectory pairs is chosen as 10 times the number of trajectories. Preference tags were then generated for these trajectory pairs in the mixed datasets. To adjust the randomness of the preferences, a steepness parameter was introduced as a scalar of the standardized reward. This configuration ensures a comprehensive dataset that can effectively support the evaluation of our methods.
@@ B.2 Tasks Descriptions
    • Challenge: The adversaries only gets reward at the timestep of catching a good agent, so recovering the reward distribution across time becomes a challenging work. Note that the original environment is a 1v3 adversary game, and we convert it into a 3-agent cooperative game for better evaluation by fixing the good agent with a MAPPO pretrained policy. This environments requires high operation precision.
@@
  • Overcooked
    • Objective: Two agents in a gridworld scoring points by repeatedly completing a three-step process: gathering ingredients, cooking, and serving dishes. Each step requires interacting with the environment at specific locations and orientations.
    • Challenge: The episode length in this environment is notably long (400 timesteps), with sparse reward signals. This creates significant challenges for training the reward model, as it may incorrectly attribute rewards to specific behaviors, resulting in inaccurate training objectives and suboptimal learning outcomes.